OpenStudy (anonymous):

Which of the following best describes the solution to the equation below? 30x + 27 = -45x - 6 A. exactly one real solution, B. no real solutions C. infinite real solutions D. exactly one real solution,

OpenStudy (anonymous):

Just solve it and you'll see.

OpenStudy (amoodarya):

30x + 27 = -45x - 6 30x + 45x = 27 - 6

OpenStudy (amoodarya):

75x=21 so x=21/75 can you go on ?